Write an equation for a parabola: Focus (4,0) , Directrix x=-4
Answer by katplus1(2) 
 Since the Directrix is horizontal, use the equation of a parabola that opens left or right.
(y-k)^2=4p(x-h)
Find the vertex: x = x coordinate of focus + directrix/2. The y coordinate will be the same as the y coordinate of the focus.
(4-4/2,0)
(0,0)
Find the distance from the focus to the vertex and from the vertex to the directrix is |p|. Subtract the x coordinate of the vertex from the x coordinate of the focus to find p.
p=4-0
p=4
Substitute in the known values for the variables into the equation
(y-k)^2=4p(x-h)
(y-0)^2=4(4)(x-0)
Simplify
y^2=16x
|
|
|
| |